# Evolution of the Inner Accretion Flow in Swift J1727.8–1613 across Intermediate States: Insights from Broadband Spectral and Timing Analysis

### v1

@@ 5 Conclusions
The broadband spectrum up to $78$ keV in the HIMS can be described with models having two different Comptonizing regions, with the electron distribution being predominantly thermal or Maxwellian. We did not notice any contribution from the non-thermal electron population in the spectral data up to 78 keV in the HIMS. The disk-corona geometry of Swift J1727.8–1613 shows strong similarity with that of GX 339–4 in the HIMS (Chand2024ApJ...972...20C).
@@
The contributions from the non-thermal electron distribution become significant above 100 keV, where a weak hard X-ray tail is observed. The shape of this tail can be satisfactorily described considering a hybrid electron distribution in the corona, where the electrons are thermally populated up to a certain energy, and then become non-thermal beyond the thermal cutoff.
@@
The low electron temperature of the hard Comptonizing region could result from the energy balance within the hot flow or the efficient cooling by synchrotron photons produced by a hybrid electron plasma, which can prevent the electrons from reaching higher temperatures.
@@
We find from the reflection spectroscopy that the accretion disk is likely to be truncated away by at least $33~r_{\rm{g}}$ from the ISCO in the HIMS, which also shows close agreement with the values estimated from modeling the disk continuum. The truncated disk scenario matches well with previous results reported for the bright LHS of other low-mass BHXRBs under the framework of dual Comptonizing regions (Zdziarski2021ApJ...909L...9Z; Banerjee2024ApJ...964..189B; Chand2024ApJ...972...20C).
@@
The source exhibits a transition to the SIMS during Obs2 from the HIMS in Obs1 with an elevated disk temperature and a relatively steeper photon index. Unlike in the HIMS, the requirement of a single Comptonizing region to describe the broadband spectrum in the SIMS suggests a significant change in the coronal geometry across the state transition. This interpretation is further supported by the disk reaching relatively closer to the ISCO and a significant increase in the QPO frequency during the SIMS. However, the lack of high-energy data prevents us from probing the existence of a hard X-ray tail that may originate from the acceleration of a non-thermal electron population.
@@
Our estimates of the black hole mass, spin, and disk inclination angle, obtained by modeling the disk continuum and reflection component together, broadly agree with the previously reported values. Within the uncertainties, these parameters are found to be consistent across the different models. Additionally, the distance to the source measured using model M5 closely matches the value reported by Mata2025A&A...693A.129M based on optical spectroscopy.
@@
The shape of the rms–energy spectra is consistent with those observed in the intermediate states of other low-mass BHXRBs. Our results indicate that the accretion disk remains relatively stable or only moderately variable, while the Comptonizing component exhibits strong variability in both spectral states. This behavior suggests that the overall source variability is primarily governed by the Comptonizing region.
6 Acknowledgments
@@
We would like to thank the anonymous reviewer for his/her comments that have helped to improve the quality of this paper. This research has made use of archival data from the NICER and NuSTAR missions, obtained from the High Energy Astrophysics Science Archive Research Center (HEASARC), provided by NASA’s Goddard Space Flight Center. It also makes use of data from the CZTI instrument onboard AstroSat, a mission of the Indian Space Research Organisation (ISRO), with data archived at the Indian Space Science Data Centre (ISSDC). This work utilized data and software from the Insight-HXMT mission, a project funded by the China National Space Administration (CNSA) and the Chinese Academy of Sciences (CAS). We also acknowledge the use of data from the MAXI mission provided by RIKEN, JAXA, and the MAXI team. SC acknowledges support from the Polish Academy of Sciences through their study visit program. AAZ acknowledges support from the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138.

### v2

@@ V Conclusions
The broadband spectrum up to $78$ keV in HIMS can be described by models with two Comptonizing regions, in which the electron distribution is predominantly thermal (Maxwellian). We did not observe any contribution from the non-thermal electron population in the HIMS spectral data up to 78 keV. The disk-corona geometry of Swift J1727.8–1613 shows substantial similarity with that of GX 339–4 in the HIMS (Chand et al., 2024).
@@
The contributions from the non-thermal electron distribution become significant above 100 keV, where a weak hard X-ray tail is observed. The shape of this tail can be satisfactorily described considering a hybrid electron distribution in the corona, where the electrons are thermally populated up to a specific energy, and then become non-thermal beyond the thermal cutoff.
@@
The low electron temperature in the hard Comptonizing region could result from an energy balance within the hot flow or from efficient cooling by synchrotron photons produced by a hybrid electron plasma, which can prevent electrons from reaching higher temperatures.
@@
We find from reflection spectroscopy that the accretion disk is likely truncated by at least $33r_{\rm{g}}$ from the ISCO in the HIMS, which also agrees closely with values estimated from modeling the disk continuum. The truncated disk scenario is consistent with previous results for the bright LHS of other low-mass BHXRBs within the framework of dual Comptonizing regions (Zdziarski et al., 2021b; Banerjee et al., 2024; Chand et al., 2024).
@@
The source transitions from HIMS in Obs1 to SIMS in Obs2, with an elevated disk temperature and a relatively steeper photon index. Unlike in HIMS, the requirement for a single Comptonizing region to describe the broadband spectrum in SIMS suggests a significant change in coronal geometry across the state transition. This interpretation is further supported by the disk’s approach to the ISCO and by a substantial increase in the QPO frequency during the SIMS. However, the lack of high-energy data prevents us from probing the existence of a hard X-ray tail that may originate from the acceleration of a non-thermal electron population.
@@
Our estimates of the black hole mass, spin, and disk inclination angle, obtained by jointly modeling the disk continuum and reflection components, broadly agree with previously reported values. Despite the uncertainties, these parameters are consistent across the models.
@@
The shape of the rms–energy spectra is consistent with those observed in the intermediate states of other low-mass BHXRBs. Our results indicate that the accretion disk remains relatively stable or only moderately variable, while the Comptonizing component exhibits substantial variability in both spectral states. This behavior suggests that the Comptonizing region primarily governs the overall source variability.
VI Acknowledgments
@@
We thank the anonymous reviewer for their comments, which have helped improve the quality of this paper. This research used archival data from the NICER and NuSTAR missions, obtained from the High Energy Astrophysics Science Archive Research Center (HEASARC) and provided by NASA’s Goddard Space Flight Center. It also utilizes data from the CZTI instrument onboard AstroSat, a mission of the Indian Space Research Organisation (ISRO), archived at the Indian Space Science Data Centre (ISSDC). This work used data and software from the Insight-HXMT mission, a project funded by the China National Space Administration (CNSA) and the Chinese Academy of Sciences (CAS). We also acknowledge the use of MAXI mission data, provided by RIKEN, JAXA, and the MAXI team. We thank N. P. S. Mithun from the CZTI team for valuable discussions on the AstroSat/CZTI background. SC thanks Hsiang-Kuang Chang for his support. SC acknowledges support from the Polish Academy of Sciences through their study visit program. AAZ acknowledges support from the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138.